Home Hardware reported beginning inventory of twenty two shovels, for a total cost of $110. The company had the following traJournal entry worksheet Record the cost of goods sold Note: Enter debits before credits. General Journal Debit Credit TransacJournal entry worksheet Record the inventory on hand based on inventory count Note: Enter debits before credits. Transaction2. Prepare the journal entries that would be recorded using a perpetual inventory system, including any book-to-physical adjuJournal entry worksheet 2 3 4 5 6 7 8 9 Record the sale of shovels. Note: Enter debits before credits. General Journal DebitJournal entry worksheet 3 4 5 6 7 8 9 Record the cost of shovels sold. Note: Enter debits before credits. General Journal DebJournal entry worksheet < 1 2 3 4 5 6 7 8 9 Record the sale of shovels. Note: Enter debits before credits. General Journal DeJournal entry worksheet Record the cost of shovels sold. Note: Enter debits before credits. General Journal Debit Credit TranJournal entry worksheet < 1 2 3 4 7 8 9 Record the purchase of shovels. Note: Enter debits before credits. General Journal DeJournal entry worksheet Record the sale of shovels. Note: Enter debits before credits. General Journal Debit Credit TransactiJournal entry worksheet < 1 2 3 4 5 6 7 8 9 Record the cost of shovels sold. Note: Enter debits before credits. Transaction GJournal entry worksheet < 1 2 3 4 5 6 7 8 9 Record the purchase of shovels. Note: Enter debits before credits. Transaction GeJournal entry worksheet < 1 2 3 4 5 6 7 8 9 Record the entry for book-to-physical adjustment, if any. Note: Enter debits befo3-a. What is the dollar amount of shrinkage that you were able to determine in requirement 1? (Enter O if you were unable t
Answer #1

1)

Periodic Inventory System
Transaction General Journal Debit Credit
Jan. 2 Accounts Receivable $72
   Sales Revenue (6 shovels * $12 per unit) $72
(To record the sales made on account)
Jan. 16 Accounts Receivable $144
   Sales Revenue (12 shovels * $12 per unit) $144
(To record the sales made on account)
Jan. 18 Purchases (7 shovels * $5 per unit) $35
   Accounts Payable $35
(To record the purchase of merchanise inventory on account)
Jan. 19 Accounts Receivable $132
   Sales Revenue (11 shovels * $12 per unit) $132
(To record the sales made on account)
Jan. 24 Purchases (12 shovels * $5 per unit) $60
   Accounts Payable $60
(To record the purchase of merchanise inventory on account)
Jan. 31 Shrinkage Loss (22 - 6 - 12 + 7 - 11 + 12 = 12 - 9 = 3 * $5) $15
   Purchases $15
(To record the shrinkage loss)
Jan. 31 Merchandise Invetory $85
   purchases ($35 + $60 - $15) $85
(To record the adjusting entry to transfer amount from purchases to Merchandising Inventory)

2)

Perpetual Inventory System
Transaction General Journal Debit Credit
Jan. 2 Accounts Receivable $72
   Sales Revenue (6 shovels * $12 per unit) $72
(To record the sales made on account)
Jan. 16 Accounts Receivable $144
   Sales Revenue (12 shovels * $12 per unit) $144
(To record the sales made on account)
Jan. 18 Merchandise Inventory (7 shovels * $5 per unit) $35
   Accounts Payable $35
(To record the purchase of merchanise inventory on account)
Jan. 19 Accounts Receivable $132
   Sales Revenue (11 shovels * $12 per unit) $132
(To record the sales made on account)
Jan. 24 Merchandise Inventory (12 shovels * $5 per unit) $60
   Accounts Payable $60
(To record the purchase of merchanise inventory on account)
Jan. 31 Shrinkage Loss (22 - 6 - 12 + 7 - 11 + 12 = 12 - 9 = 3 * $5) $15
   Merchandise Inventoy $15
(To record the shrinkage loss)

The dollar amount of shrinkage is $15 as per requirement 1.

The dollar amount of shrinkage is $15 as per requirement 2.